# Inference-Time Scaling of Discrete Diffusion Models via Importance Weighting and Optimal Proposal Design

### v3

@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020) have achieved remarkable success across various domains, from image synthesis (Rombach et al., 2022; Esser et al., 2024) to scientific applications (Hoogeboom et al., 2022; Watson et al., 2023). Recently, advances in discrete diffusion models (Austin et al., 2021; Sahoo et al., 2024; Shi et al., 2024) have established them as a powerful approach for modelling discrete data, notably in tasks such as language modelling (Nie et al., 2025; Zhang et al., 2025) and code generation (Gat et al., 2024; Gong et al., 2025).
Despite their impressive capabilities, pretrained diffusion models often need to generate samples that meet specific downstream constraints. For example, text-to-image generation requires images aligned with human preferences (Black et al., 2023; Fan et al., 2023; Uehara et al., 2024), while protein generation demands stability or desired binding affinity (Verkuil et al., 2022; Uehara et al., 2025b). To address this challenge, existing approaches mainly fall into two categories: i) fine-tuning and ii) guidance methods. Fine-tuning methods, including techniques such as steering (Rector-Brooks et al., 2024), reinforcement learning (Zekri & Boullé, 2025), and direct backpropagation (Wang et al., 2024), have demonstrated promising results. Nevertheless, these methods often suffer from reward over-optimisation, which can compromise sample quality and diversity. On the other hand, guidance and sampling methods (Li et al., 2024; Gruver et al., 2023; Nisonoff et al., 2024; Guo et al., 2024; Uehara et al., 2025a) provide training-free alternatives that are easier to deploy, but they often suffer from reward under-optimisation. This limits their ability to enforce correct alignment, resulting in outputs that may not fully meet complex constraints.
@@ 2.2 Importance Sampling and Sequential Monte Carlo
${\mathbb{E}}_{\pi(x_{t})}[\delta(x_{t})]={\mathbb{E}}_{q(x_{t:T})}\left[\frac{\pi(x_{t:T})}{q(x_{t:T})}\delta(x_{t})\right]\approx\frac{1}{N}\sum_{i=1}^{N}\frac{\pi(x_{t:T}^{(i)})}{q(x_{t:T}^{(i)})}\delta(x_{t}),\quad x_{t:T}^{(i)}\sim q(x_{t:T}).$ (4)
@@
$w_{t-1}(x_{t-1:T})\triangleq\frac{\pi(x_{t-1:T})}{q(x_{t-1:T})}=\frac{\pi(x_{t-1}|x_{t:T})\pi(x_{t:T})}{q(x_{t-1}|x_{t:T})q(x_{t:T})}=\frac{\pi(x_{t-1})}{\pi(x_{t})}\frac{\gamma(x_{t}|x_{t-1})}{q(x_{t-1}|x_{t})}w_{t}(x_{t:T}),$ (5)
where we leverage the Markovian assumption that $\pi(x_{t:T})=\pi(x_{t})\prod_{k=t}^{T-1}\gamma(x_{t+1}|x_{t})$ for arbitrary forward kernel $\gamma$ and thus $\pi(x_{t-1}|x_{t})=\pi(x_{t-1})\gamma(x_{t}|x_{t-1})/\pi(x_{t})$. The recursion of importance weight underlies the iterative procedure of SMC. Concretely, The procedure initialises begins by $N$ particles $x_{T}^{(i)}\sim q(x_{T})$ with weights $w_{T}^{(i)}\leftarrow\pi(x_{T}^{(i)})/q(x_{T}^{(i)})$. For each step $t=T,\dots,1$ and particles $i=1,\dots,N$, SMC proceeds as follows: i) resample ancestor according to the weights $\{w_{t}^{(i)}\}_{i=1}^{N}$; ii) propagate new particles via $x_{t-1}^{(i)}\sim q(x_{t-1}|x_{t})$; and iii) updating the weights as $w_{t-1}^{(i)}\leftarrow[\pi(x_{t-1})\pi(x_{t}|x_{t-1})]/[\pi(x_{t})q(x_{t-1}|x_{t})]$. The resulting collection of weighted particles provides an asymptotically consistent approximation of the intermediate target distribution $\pi(x_{t})$.
@@ 3 Sequence Monte Carlo for Discrete Diffusion Models
Given a pretrained discrete diffusion model $p_{\theta}(x_{t})$, we consider sampling from modified target distributions that enable inference-time control. These targets include: i) product distributions, a general form underlying classifier free guidance (Ho & Salimans, 2022), defined as $\pi\propto p_{\theta_{1}}^{\alpha}(x_{t})p_{\theta_{2}}^{\beta}(x_{t})$; and ii) reward-tilting distributions, expressed as $\pi\propto p_{\theta}(x_{t})\exp(r(x_{t}))$. In the following section, we introduce how to construct tractable importance weights by carefully selecting forward kernels $\gamma(x_{t}|x_{t-1})$, and show their connection to existing SMC formulations for continuous-time discrete diffusion models. We then discuss the choice of proposal distributions, which play a central role in balancing variance reduction with computational efficiency.
@@ 3.1 Importance Weight: Tractability with Pretrained Diffusion Models
To perform SMC, one must evaluate the importance weight from Equation˜5 at each step $t$. While the forward kernel $\gamma(x_{t}|x_{t-1})$ and the proposal $q(x_{t-1}|x_{t})$ can be chosen flexibly, the ratio of intermediate targets $\frac{\pi(x_{t-1})}{\pi(x_{t})}$ is generally intractable in diffusion models. With a well-trained diffusion model $p_{\theta}$, however, this ratio can be approximated via detailed balance $\frac{p_{\theta}(x_{t-1})}{p_{\theta}(x_{t})}\approx\frac{p_{\theta}(x_{t-1}|x_{t})}{p(x_{t}|x_{t-1})}$, where $p(x_{t}|x_{t-1})$ denotes the forward noising process and $p_{\theta}(x_{t-1}|x_{t})$ is learned reverse counterpart. Under this approximation, the importance weight for the product target takes the form
@@
$\text{reward-tilting:}\ \frac{\exp(r(x_{s}))}{\exp(r(x_{t}))}\frac{p_{\theta}(x_{t-1}|x_{t})}{p(x_{t}|x_{t-1})}\frac{\gamma(x_{t}|x_{t-1})}{q(x_{t-1}|x_{t})}=\frac{\exp(r(x_{t-1}))}{\exp(r(x_{t}))}\frac{p_{\theta}(x_{t-1}|x_{t})}{q(x_{t-1}|x_{t})}$ (7)
@@
Let $R_{t}$ be the rate matrix generating the forward transition kernel $\gamma(x_{t}|x_{t-\Delta t})$, and $\hat{R}_{t}$ be its counterpart associated with the backward proposal kernel $q(x_{t-\Delta t}|x_{t})$, where $\Delta t\to 0$ is the infinitesimal time increment. Then, the importance weight at time $t$ is given by $w_{t}=\int_{1}^{t}-\partial_{s}\log\pi(x_{s})+\sum_{y_{s}}R_{s}(x_{s},y_{s})\frac{\pi(y_{s})}{\pi(x_{s})}\mathop{}\!\mathrm{d}s$.
@@ 3.2 Choices of Proposal: the Way to Reduce Variance
Building on Proposition˜2, one readily verifies that the optimal proposal distribution for the case of product target is tractable (see the remark in Section˜A.1 for discussion), under the choices of forward kernels $\gamma\propto p_{1}^{\alpha}(x_{t}|x_{t-1})p_{2}^{\beta}(x_{t}|x_{t-1})$. In contrast, for the reward-tilting, the locally optimal proposal takes the form $q\propto\exp(r(x_{t-1}))p_{\theta}(x_{t-1}|x_{t})$, which is generally computationally infeasible due to the inaccessibility of the corresponding normalising constant. Consequently, practical implementations must resort to approximations that balance variance reduction with computational feasibility. In what follows, we introduce two approximation strategies tailored to the reward-tilting setting: i) a gradient-based method to achieve first-order approximation, and ii) a neural proposal trained to minimise the log-variance of the importance weight.
@@ 3.2.1 Approximated Optimal Proposal via First-Order Approximation
This approximation improves computational efficiency by requiring the reward function $r$ to be evaluated and differentiated only once at $x_{t}$, instead of repeatedly across all states. Nevertheless, it assumes differentiable rewards and remains costly when the reward model is large. Motivated by Richter et al. (2020); Richter & Berner (2023) and the amortisation technique in variational inference (Dayan et al., 1995; Kingma & Welling, 2013), we propose learning an amortised network that approximates the optimal proposal. This reduces computation to a single network evaluation, thereby significantly enhancing the efficiency of SMC.
@@ 3.3 Sequential Monte Carlo Recipe: Practical Implementation
However, categorical sampling from $p_{\theta}$ renders $\hat{r}(x_{t})$ non-differentiable w.r.t. $x_{t}$. To resolve it, we employ the reparameterisation trick with Gumbel-Softmax (Jang et al., 2016) to enable differentiability, thereby making the approximated proposal applicable as in Equation˜8. For completeness, Section˜C.2 provides a further discussion of the computation of importance weights in Equation˜7 under low-confidence sampling, where the ratio $\frac{p_{\theta}(x_{t-1}|x_{t})}{q(x_{t-1}|x_{t})}$ is not explicitly tractable. This extension ensures that the proposed SMC algorithm remains suitable for recent state-of-the-art discrete diffusion models for language modelling (Nie et al., 2025) and text-to-image generation (Bai et al., 2024), where low-confidence sampling (Chang et al., 2022) is commonly used.
@@ 4.2 Language Modelling
Following Han et al. (2022), we generate sequences of length 100 with 100 denoising steps condition on the given starting prompts, and report results averaged over 300 independent runs corresponding to 15 prompts with 20 generations per prompt. The results are summarised in Table˜1, with an extended version provided in Table˜4. We observe that SMC with proposals closer to the optimal achieves better performance on the toxicity metrics, reflecting stronger alignment with the reward model. Among the non-SMC baselines, $\text{Prop}_{\text{amot}}$ yields the best performance, highlighting the effectiveness of the log-variance minimisation objective. To further assess its effect, we plot the training dynamics in Figure˜7, which shows the reward steadily improving as training progresses. Notably, although the learned proposal sacrifices a small degree of performance on perplexity and diversity, we demonstrate in Figure˜16 that it consistently generates coherent and semantically meaningful sequences, indicating that alignment improvements need not come at the expense of sample quality.
@@ 4.3 Biology Design
As shown in Figure˜4, the performance improves consistently with an increasing number of particles, suggesting that SMC benefits from a larger particle set by providing a more accurate approximation of the target distribution. Compared to $\text{SMC}_{\text{base}}$, $\text{SMC}_{\text{amot}}$ achieves higher Pred-Activity and ATAC-Acc, while performing slightly worse on the other three metrics. This can be attributed to the more mode-seeking behaviour of their proposals, which emphasises high-probability regions at the expense of overall diversity. Nonetheless, we observe that $\text{SMC}_{\text{amot}}$ with larger particle sets attains more favourable overall performance, indicating that a learnable amortised proposal can effectively leverage the flexibility of SMC to balance quality and diversity. In Section˜D.2.3, we present additional comparisons with baseline methods, further demonstrating the effectiveness of our approach.
@@ 4.4 Image Generation
Improving CFG with MaskGit (Chang et al., 2022). Given a pretrained diffusion model $p_{\theta}$, classifier-free guidance (CFG) generates samples according to $p_{\theta}(x_{t-1}|x_{t},c)^{\alpha}p_{\theta}(x_{t-1}|x_{t})^{1-\alpha}$, where $\alpha$ is the CFG coefficient. CFG has been shown to enhance sample quality substantially (Ho & Salimans, 2022). By incorporating the importance weight defined in Equation˜6, we can further improve CFG within the proposed SMC framework.
@@
Specifically, we perform experiments with MaskGit (Chang et al., 2022) trained on ImageNet256 (Deng et al., 2009). To ensure tractable importance weight computation, we adopt the ReMDM sampling scheme (Wang et al., 2025) instead of the low-confidence sampling strategy from (Chang et al., 2022) (see Tables˜10, 10 and 10 for a comparison). The result is presented in Table˜2. It shows that with fewer denoising steps, increasing the number of particles leads to a substantial improvement in sample quality, as measured by FID and Inception Score (IS) with $50,000$ generated images, thereby demonstrating the effectiveness of the proposed SMC method. However, as the number of denoising steps increases, the benefit of using more particles diminishes. This can be attributed to the role of SMC in correcting sampling inaccuracies: with sufficient denoising steps, the sampling process itself becomes accurate enough, leaving limited room for further improvement through additional particles. In addition, we report results with different CFG coefficients in Tables˜10 and 10. Interestingly, for larger CFG coefficients, increasing the particle count tends to decrease FID while increasing IS. This behaviour is expected, since stronger CFG reduces sample diversity. With more accurate sampling under SMC, this reduction in diversity becomes more apparent, leading to lower FID but higher IS.
@@
The results are shown in Figure˜5. We observe that performance consistently improves with an increasing number of particles, and $\text{SMC}_{\text{amot}}$ outperforms all other methods, which highlights the benefit of the proposed SMC framework. In Figure˜6, we visualise the alignment dynamics for the HPSv2 task, showing that the generated images progressively align more faithfully with the prompts, thereby validating the effectiveness of the proposed log-variance minimisation objective. Furthermore, Figures˜7, 7 and 7 present the convergence of the reward during training, and further qualitative examples in Section˜D.2.5 collectively reinforce the validity of our approach.

### v4

@@ 1 Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020) have achieved remarkable success across various domains, from image synthesis (Rombach et al., 2022; Esser et al., 2024) to scientific applications (Hoogeboom et al., 2022; Watson et al., 2023). Recently, advances in discrete diffusion models (Austin et al., 2021; Sahoo et al., 2024; Shi et al., 2024) have established them as a powerful approach for modelling discrete data, notably in tasks such as language modelling (Nie et al., 2025; Zhang et al., 2025a) and code generation (Gat et al., 2024; Gong et al., 2025).
Despite their impressive capabilities, pretrained diffusion models often need to generate samples that meet specific downstream constraints. For example, text-to-image generation requires images aligned with human preferences (Black et al., 2023; Fan et al., 2023; Uehara et al., 2024), while protein generation demands stability or desired binding affinity (Verkuil et al., 2022; Uehara et al., 2025b). To address this challenge, existing approaches mainly fall into two categories: i) fine-tuning and ii) guidance methods. Fine-tuning methods, including techniques such as steering (Rector-Brooks et al., 2024), reinforcement learning (Zekri and Boullé, 2025), and direct backpropagation (Wang et al., 2024), have demonstrated promising results. Nevertheless, these methods often suffer from reward over-optimisation, which can compromise sample quality and diversity. On the other hand, guidance and sampling methods (Li et al., 2024; Gruver et al., 2023; Nisonoff et al., 2024; Guo et al., 2024; Uehara et al., 2025a) provide training-free alternatives that are easier to deploy, but they often suffer from reward under-optimisation. This limits their ability to enforce correct alignment, resulting in outputs that may not fully meet complex constraints.
@@ 2.2 Importance Sampling and Sequential Monte Carlo
${\mathbb{E}}_{\pi(x_{t})}[\delta(x_{t})]={\mathbb{E}}_{q(x_{t:T})}\left[\frac{\pi(x_{t:T})}{q(x_{t:T})}\delta(x_{t})\right]\approx\frac{1}{N}\sum_{i=1}^{N}\frac{\pi(x_{t:T}^{(i)})}{q(x_{t:T}^{(i)})}\delta(x_{t}^{(i)}),\quad x_{t:T}^{(i)}\sim q(x_{t:T}).$ (4)
@@
$w_{t-1}(x_{t-1:T}^{(i)})\triangleq\frac{\pi(x_{t-1:T}^{(i)})}{q(x_{t-1:T}^{(i)})}=\frac{\pi(x_{t-1}^{(i)}|x_{t:T}^{(i)})\pi(x_{t:T}^{(i)})}{q(x_{t-1}^{(i)}|x_{t:T}^{(i)})q(x_{t:T}^{(i)})}=\frac{\pi(x_{t-1}^{(i)})}{\pi(x_{t}^{(i)})}\frac{\gamma(x_{t}^{(i)}|x_{t-1}^{(i)})}{q(x_{t-1}^{(i)}|x_{t}^{(i)})}w_{t}(x_{t:T}^{(i)}),$ (5)
where we leverage the Markovian assumption that $\pi(x_{t:T}^{(i)})=\pi(x_{t}^{(i)})\prod_{k=t}^{T-1}\gamma(x_{t+1}^{(i)}|x_{t}^{(i)})$ for arbitrary forward kernel $\gamma$ and thus $\pi(x_{t-1}^{(i)}|x_{t}^{(i)})=\pi(x_{t-1}^{(i)})\gamma(x_{t}^{(i)}|x_{t-1}^{(i)})/\pi(x_{t}^{(i)})$. The recursion of importance weight underlies the iterative procedure of SMC. Concretely, The procedure initialises begins by $N$ particles $x_{T}^{(i)}\sim q(x_{T})$ with weights $w_{T}^{(i)}\leftarrow\pi(x_{T}^{(i)})/q(x_{T}^{(i)})$. For each step $t=T,\dots,1$ and particles $i=1,\dots,N$, SMC proceeds as follows: i) resample ancestor according to the weights $\{w_{t}^{(i)}\}_{i=1}^{N}$; ii) propagate new particles via $x_{t-1}^{(i)}\sim q(x_{t-1}|x_{t})$; and iii) updating the weights as $w_{t-1}^{(i)}\!\leftarrow\![\pi(x_{t-1}^{(i)})\pi(x_{t}^{(i)}|x_{t-1}^{(i)})]/[\pi(x_{t}^{(i)})q(x_{t-1}^{(i)}|x_{t}^{(i)})]$. The resulting collection of weighted particles provides an asymptotically consistent approximation of the intermediate target distribution $\pi(x_{t})$.
@@ 3 Sequence Monte Carlo for Discrete Diffusion Models
Given a pretrained discrete diffusion model $p_{\theta}(x_{t})$, we consider sampling from modified target distributions that enable inference-time control. These targets include: i) product distributions, a general form underlying classifier free guidance (Ho and Salimans, 2022), defined as $\pi(x_{t})\propto p_{\theta_{1}}^{\alpha}(x_{t})p_{\theta_{2}}^{\beta}(x_{t})$; and ii) reward-tilting distributions, expressed as $\pi(x_{t})\propto p_{\theta}(x_{t})\exp(r(x_{t}))$. In the following section, we introduce how to construct tractable importance weights by carefully selecting forward kernels $\gamma(x_{t}|x_{t-1})$, and show their connection to existing SMC formulations for continuous-time discrete diffusion models. We then discuss the choice of proposal distributions, which play a central role in balancing variance reduction with computational efficiency.
@@ 3.1 Importance Weight: Tractability with Pretrained Diffusion Models
To perform SMC, one must evaluate the importance weight from Equation˜5 at each step $t$. While the forward kernel $\gamma(x_{t}|x_{t-1})$ and the proposal $q(x_{t-1}|x_{t})$ can be chosen flexibly, the ratio of intermediate targets $\frac{\pi(x_{t-1})}{\pi(x_{t})}$ is generally intractable in diffusion models. With a well-trained diffusion model $p_{\theta}$, however, this ratio can be approximated via detailed balance $\frac{p_{\theta}(x_{t-1})}{p_{\theta}(x_{t})}\approx\frac{p_{\theta}(x_{t-1}|x_{t})}{p(x_{t}|x_{t-1})}$, where $p(x_{t}|x_{t-1})$ denotes the forward noising process and $p_{\theta}(x_{t-1}|x_{t})$ is the learned reverse counterpart. Under this approximation, the importance weight for the product target takes the form
@@
$\text{reward-tilting:}\ \frac{\exp(r(x_{t-1}))}{\exp(r(x_{t}))}\frac{p_{\theta}(x_{t-1}|x_{t})}{p(x_{t}|x_{t-1})}\frac{\gamma(x_{t}|x_{t-1})}{q(x_{t-1}|x_{t})}=\frac{\exp(r(x_{t-1}))}{\exp(r(x_{t}))}\frac{p_{\theta}(x_{t-1}|x_{t})}{q(x_{t-1}|x_{t})}$ (7)
@@
Let $R_{t}$ be the rate matrix generating the forward transition kernel $\gamma(x_{t}|x_{t-\Delta t})$, and $\hat{R}_{t}$ be its counterpart associated with the backward proposal kernel $q(x_{t-\Delta t}|x_{t})$, where $\Delta t\to 0$ is the infinitesimal time increment. Then, the importance weight at time $t$ is given by $w_{t}=\int_{1}^{t}-\partial_{s}\log\pi(x_{s})+\sum_{y_{s}}R_{s}(x_{s},y_{s})\frac{\pi(y_{s})}{\pi(x_{s})}\mathop{}\!\mathrm{d}s$, if the forward kernel $\gamma$ is chosen such that the rate matrices satisfy detailed balance $\hat{R}_{t}(x_{t},y_{t})\pi(x_{t})={R}_{t}(y_{t},x_{t})\pi(y_{t})$.
@@ 3.2 Choices of Proposal: the Way to Reduce Variance
Building on Proposition˜2, one readily verifies that the optimal proposal distribution for the case of product target is tractable (see the remark in Section˜A.2 for discussion), under the choices of forward kernels $\gamma\propto p_{1}^{\alpha}(x_{t}|x_{t-1})p_{2}^{\beta}(x_{t}|x_{t-1})$. In contrast, for the reward-tilting, the locally optimal proposal takes the form $q\propto\exp(r(x_{t-1}))p_{\theta}(x_{t-1}|x_{t})$, which is generally computationally infeasible due to the inaccessibility of the corresponding normalising constant. Consequently, practical implementations must resort to approximations that balance variance reduction with computational feasibility. In what follows, we introduce two approximation strategies tailored to the reward-tilting setting: i) a gradient-based method to achieve first-order approximation, and ii) a neural proposal trained to minimise the log-variance of the importance weight.
@@ 3.2.1 Approximated Optimal Proposal via First-Order Approximation
This approximation improves computational efficiency by requiring the reward function $r$ to be evaluated and differentiated only once at $x_{t}$, instead of repeatedly across all states. Nevertheless, it assumes differentiable rewards and remains costly when the reward model is large. Motivated by Richter et al. (2020); Richter and Berner (2023) and the amortisation technique in variational inference (Dayan et al., 1995; Kingma and Welling, 2013), we propose learning an amortised network that approximates the optimal proposal, resulting in a transition kernel that directly transports between successive intermediate targets as in Matthews et al. (2022). This reduces computation to a single network evaluation, thereby significantly enhancing the efficiency of SMC.
@@ 3.3 Sequential Monte Carlo Recipe: Practical Implementation
However, categorical sampling from $p_{\theta}$ renders $\hat{r}(x_{t})$ non-differentiable w.r.t. $x_{t}$. To resolve it, we employ the reparameterisation trick with Gumbel-Softmax (Jang et al., 2016) to enable differentiability (see Section˜C.2 for details), thereby making the approximated proposal applicable as in Equation˜8. For completeness, Section˜C.3 provides a further discussion of the computation of importance weights in Equation˜7 under low-confidence sampling, where the ratio $\frac{p_{\theta}(x_{t-1}|x_{t})}{q(x_{t-1}|x_{t})}$ is not explicitly tractable. This extension ensures that the proposed SMC algorithm remains suitable for recent state-of-the-art discrete diffusion models for language modelling (Nie et al., 2025) and text-to-image generation (Bai et al., 2024), where low-confidence sampling (Chang et al., 2022) is commonly used.
@@ 4.2 Language Modelling
Following Han et al. (2022), we generate sequences of length 100 with 100 denoising steps condition on the given starting prompts, and report results averaged over 300 independent runs corresponding to 15 prompts with 20 generations per prompt. The results are summarised in Table˜1, with an extended version provided in Table˜4. We observe that SMC with proposals closer to the optimal achieves better performance on the toxicity metrics, reflecting stronger alignment with the reward model. Among the non-SMC baselines, $\text{Prop}_{\text{amot}}$ yields the best performance, highlighting the effectiveness of the log-variance minimisation objective. To further assess its effect, we plot the training dynamics in Figure˜7, which shows the reward steadily improving as training progresses. Notably, although the learned proposal sacrifices a small degree of performance on perplexity and diversity, we demonstrate in Figure˜19 that it consistently generates coherent and semantically meaningful sequences, indicating that alignment improvements need not come at the expense of sample quality.
@@ 4.3 Biology Design
As shown in Figure˜4, the performance improves consistently with an increasing number of particles, suggesting that SMC benefits from a larger particle set by providing a more accurate approximation of the target distribution. Compared to $\text{SMC}_{\text{base}}$, $\text{SMC}_{\text{amot}}$ achieves higher Pred-Activity and ATAC-Acc, while performing slightly worse on the other three metrics. This can be attributed to the more mode-seeking behaviour of their proposals, which emphasises high-probability regions at the expense of overall diversity. Nonetheless, we observe that $\text{SMC}_{\text{amot}}$ with larger particle sets attains more favourable overall performance, indicating that a learnable amortised proposal can effectively leverage the flexibility of SMC to balance quality and diversity. In Section˜D.2.4, we present additional comparisons with baseline methods, further demonstrating the effectiveness of our approach.
@@ 4.4 Image Generation
Improving CFG with MaskGit (Chang et al., 2022). Given a pretrained diffusion model $p_{\theta}$, classifier-free guidance (CFG) generates samples according to $p_{\theta}(x_{t-1}|x_{t},c)^{\alpha}p_{\theta}(x_{t-1}|x_{t})^{1-\alpha}$, where $\alpha$ is the CFG coefficient. CFG has been shown to enhance sample quality substantially (Ho and Salimans, 2022). By incorporating the importance weight defined in Equation˜6, we can further improve CFG within the proposed SMC framework.
@@
Specifically, we perform experiments with MaskGit (Chang et al., 2022) trained on ImageNet256 (Deng et al., 2009). To ensure tractable importance weight computation, we adopt the ReMDM sampling scheme (Wang et al., 2025) instead of the low-confidence sampling strategy from (Chang et al., 2022) (see Tables˜11, 11 and 11 for a comparison). The result is presented in Table˜2. It shows that with fewer denoising steps, increasing the number of particles leads to a substantial improvement in sample quality, as measured by FID and Inception Score (IS) with $50,000$ generated images, thereby demonstrating the effectiveness of the proposed SMC method. However, as the number of denoising steps increases, the benefit of using more particles diminishes. This can be attributed to the role of SMC in correcting sampling inaccuracies: with sufficient denoising steps, the sampling process itself becomes accurate enough, leaving limited room for further improvement through additional particles. In addition, we report results with different CFG coefficients in Tables˜11 and 11. Interestingly, for larger CFG coefficients, increasing the particle count tends to decrease FID while increasing IS. This behaviour is expected, since stronger CFG reduces sample diversity. With more accurate sampling under SMC, this reduction in diversity becomes more apparent, leading to lower FID but higher IS.
@@
The results are shown in Figure˜5. We observe that performance consistently improves with an increasing number of particles, and $\text{SMC}_{\text{amot}}$ outperforms all other methods, which highlights the benefit of the proposed SMC framework. In Figure˜6, we visualise the alignment dynamics for the HPSv2 task, showing that the generated images progressively align more faithfully with the prompts, thereby validating the effectiveness of the proposed log-variance minimisation objective. Furthermore, Figures˜7, 7 and 7 present the convergence of the reward during training, and further qualitative examples in Section˜D.2.7 collectively reinforce the validity of our approach.